OpenStudy (anonymous):

Simplify 6 - 4[x + 1 - (y + 3)] + 5y

OpenStudy (anonymous):

does it give you what x equals?

OpenStudy (anonymous):

no @kaitlin_nikole

OpenStudy (anonymous):

Simplify 6 - 4[x + 1 - (y + 3)] + 5y

OpenStudy (anonymous):

Change the signs inside the parenthesis -y-3 6 - 4[x + 1 - y - 3] + 5y add the 1 and 3 together 6 - 4[x - y - 2] + 5y Multiply by 4 - 4[x - y - 2] = -4x-4y+8 That gives you... 6 -4x+4y+8 + 5y Add your y's and whole numbers 9y-4x+14

OpenStudy (anonymous):

Amir evaluated 5x - (x + 3)2 for x = 2, as shown. Describe Amir's mistake and find the correct answer. 5(2) - (2 + 3)2 = 10 + (-5)2 = 10 + 25 = 35